So this is for my really hard level 6 homework 1) 33-8t=15 and number
2)7=4u=70-3u

Answer by ewatrrr(24785)

1) 33 - 8t = 15
33 = 15 + 8t
18 = 8t
18/8 = t = 2 1/4
2)too many = signs
IF
7 +4u = 80-3u
7u = 73
u = 73/7 = 10 3/7